# Enhancing key rates of QKD protocol by Coincidence Detection

### v1

@@ IV.4 Postprocessing
(b) Simulated key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol.
(c) Experimental secure key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol. Lines drawn are for the aid of the eye.
(d) Key rates per pulse as a function of distance (in km) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) DSCD protocol. The plots are in log scale
Figure 3: (a) shows a comparison of secure key rates obtained using the traditional analytical method and the proposed numerical method. (b),(c) and (d) shows the comparison of secure key rates of the four different protocols using proposed analytical method.
@@ V Results and discussion
We compare the traditional Decoy state method with the proposed SDP-based method in calculating lower bounds on the secure key rates given in Eq  (11). This comparison is performed for various values of the mean photon number as shown in Figure. (2(a)). This comparison is done by performing a simulation of the decoy-based BB84 protocol for various mean photon numbers. In the simulation, the mean photon number of the decoy is kept constant with $\nu=0.1$. It is seen that our proposed method gives tighter lower bounds than the traditional analytic method by $27\%$ at the optimal mean photon number ($\mu=0.8$).
@@
We employ our SDP-based method to analyse the key rates for the four protocols viz. BB84, decoy state protocol, coincidence detection (CD) protocol, and decoy state coincidence detection (DSCD) protocol. We use the SDP technique to find the bounds on the yields and errors to compute the bounds on key rates using Eqs. (11) and (5). The respective key rates are plotted as a function of the mean photon number ($\mu$) in Figure. (2(b)). It is established that when coincidences are taken into account, the key rate is enhanced by a substantial amount. The coincidence detection protocol helps to achieve a higher optimal mean photon number, i.e. $\mu=1.1$, as compared to using the traditional decoy-state method without coincidences, i.e. $\mu=0.8$. Comparing the optimal key rates of decoy-state protocol and DSCD protocol, we note an improvement of $\approx 64\%$.
@@
Using the experimental setup described in the previous Section, the four protocols are executed. The protocols are systematically repeated for several signal intensities, each corresponding to distinct mean photon numbers. The experimentally acquired gain ($Q_{\mu}$) and QBER ($E_{\mu}$) are then used to estimate the optimal bounds on the secure key rate as discussed in Section. III. We have used the timestamps for recording the coincidences in addition to the single detections. The data were taken at various intensities of the source. These two cases enabled us to study four variants for our QKD protocol: Standard BB84, BB84 with decoy, BB84 with coincidences, BB84 with decoy and coincidences. Optimal key rates were calculated for all four cases at different values of mean photon number ($\mu$). The key rates for all the cases as a function of the mean photon number can be seen in Figure. (2(c)). It is also observed that our experiment reproduces the theoretical predictions of improvement in secure key rates. We obtained a maximum key rate of $0.17$ per pulse for the DSCD protocol.
@@
We have performed a simulation of the expected secure key rate for the four protocols as a function of distance in optical fiber. The simulated key rates are plotted in Figure. (2(d)). In this simulation we have considered the mean photon number of the signal and decoy to be $\mu=0.9$ and $\nu=0.1$, respectively. It is evident that the DSCD protocol consistently gives higher key rates than other protocols across varying distances, showcasing its effectiveness in extending communication ranges. This consistent performance highlights the DSCD protocol’s reliability and strength in addressing challenges related to communication distance.
@@ VI Conclusion
In this work, we have proposed the decoy state coincidence detection (DSCD) protocol, where we have integrated two approaches to harness their collective strength. Monitoring coincidences in decoy state protocol allows us to detect the most practical, sophisticated PNS attacks. With an assurance of no PNS attack, we showed, including contributions from two-photon statistics that we can achieve higher key rates.
@@ Acknowledgement
The authors acknowledge the partial funding support from DST through the QuST program. The authors are also thankful to group members of the QST lab for their valuable input, with a special thanks to Dr. Satyajeet Patil and Mr. Vardaan Mongia. RB acknowledges support from Quantum Communications Hub of the UK Engineering and Physical Sciences Research Council (EPSRC). RB thanks Mr. Lewis Wooltorton, Dr. Shashaank Ranu and Prof. Roger Coleck for insightful discussions. AB thanks Prof. Marco Lucamarini and Dr. Rupesh Kumar for insightful discussions.
@@ Appendix B Analytic bounds for key-rates
$\begin{split}R&\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\mathrm{bin}}(E_{\mu})+Q_{1}\left(1-H_{\mathrm{bin}}(e_{1})\right)\Big{\}},\end{split}$ (21)
@@ Appendix C Numerically computing key rates
Here we present an algorithm to compute tight lower bounds for the optimization problem (6).
@@ C.1 Converting into a finite optimization problem
As the objective function of problem (6) only consists of $2$ terms with $4$ variables, removing infinite parameters from constraints suffices. Consider the following lower bound on the constraint for $Q_{\nu_{d}}$ in the optimization problem:
@@
allows us to obtain a relaxation corresponding to the constraint $Q_{\mu}e^{\mu}=\sum_{k=0}^{\infty}Y_{k}\mu^{k}/k!$ in optimization problem (6):
@@
In other words, we have formed a relaxation of the constraint given by an infinite sum of infinitely many variables through a constraint that involves a finite sum of finitely many variables. We use similar arguments to derive the following relaxations of the constraints for $E_{\mu}Q_{\mu},Q_{\nu_{i}},E_{\nu_{i}}Q_{\nu_{i}}$ in optimization problem (6):
@@
$r_{n}:=\inf$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$ (26)
@@
$\forall\lambda\in S:\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@
Consider the optimization problems for $r_{n}$ and $r$. From the preceding discussion, it is evident that the constraints in problem (26) serve as relaxations for the constraints in problem (6). Consequently, the feasible set of optimization problem (6) is a subset of that of optimization problem (26). Moreover, the optimization problems designed to compute both $r$ and $r_{n}$ share identical objective functions. Thus, $r_{n}\leq r$ holds for all $n\in\mathbb{R}$.
@@ C.2 Partitioning the domain
$k_{\mathcal{P}}\leq k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})=\tilde{f}(\mathbf{x}_{i},\mathbf{y}_{i})\leq\tilde{f}_{i}(\mathbf{x}^{*},\mathbf{y}^{*})\leq f(\mathbf{x}^{*},\mathbf{y}^{*})=k.$ (30)
@@
$k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}}):=\min$ $f_{i}(\mathbf{x},\mathbf{y})$ (31)
@@
$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$
@@
$k=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})-\epsilon\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}-\epsilon\geq k_{\mathcal{P}}-\epsilon.$
@@ C.3 Algorithm for computing key-rates
$\xi^{\max}_{k,i}=\begin{cases}\max\{\Phi((2a_{k}^{(i)}-1)^{k}),\Phi((2b_{k}^{(i)}-1)^{k})\}&\text{if}\quad 0\notin[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]\\ \Phi(0)=1&\text{otherwise}\end{cases}.$ (32)
@@
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@ Appendix D Speeding up solving the optimization problem
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@ Appendix E Rates in the presence of statistical errors
$\mathfrak{A}:=\left(\bigcup_{d}\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\}\right)\cup\left(\bigcup_{d}\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\}\right)\cup\mathfrak{A}_{1}.$ (46)
@@
$\mathbb{P}(\mathfrak{A})\leq\sum_{d}\left(\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})+\mathbb{P}(\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\right)+\mathbb{P}(\mathbb{A}_{1}).$ (47)
@@
$\mathbb{P}(\mathfrak{A})\leq 2\sum_{s=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{i}}(\delta_{\nu_{d}}^{E})^{2}}\right)+2\epsilon_{\mathrm{stat}}.$ (50)

### v2

@@ IV.4 Postprocessing
(b) Simulated key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol.
(c) Experimental secure key rates per pulse as a function of mean photon number ($\mu$) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol. Lines drawn are for the aid of the eye.
(d) Key rates per pulse as a function of distance (in km) for four protocols: (i) BB84 Protocol, (ii) CD protocol, (iii) Decoy state protocol and (iv) EPCD protocol. The plots are in log scale
Figure 5: (a) shows a comparison of secure key rates obtained using the traditional analytical method and the proposed numerical method. (b),(c) and (d) show the comparison of secure key rates of the four different protocols using the proposed analytical method.
@@ V Results and discussion
Increasing the key rate in realistic scenarios remains a challenge as alternative approaches typically demand additional hardware components or intricate optimization techniques for key extraction. We proposed a solution minimizing the resource requirement on both the software and hardware side. We performed a comparative study of key rates across different protocols. We then evaluated our key rate optimization method against existing ones, with detailed results presented in subsequent paragraphs.
@@
We compare the traditional Decoy state method with the proposed SDP-based method in calculating lower bounds on the secure key rates given in Eq  (III). This comparison is performed for various values of the mean photon number as shown in Figure. (5(a)), which is done by performing a simulation of the decoy-based BB84 protocol for various mean photon numbers. In the simulation, the mean photon number of the decoy is kept constant with $\nu=0.1$. It is seen that our proposed method gives tighter lower bounds than the traditional analytic method by $27\%$ at the optimal mean photon number ($\mu=0.8$).
@@
We employ our SDP-based method to analyse the key rates for the four protocols viz. BB84, decoy state protocol, coincidence detection (CD) protocol, and entrapped pulse coincidence detection (EPCD) protocol. We use the SDP technique to find the bounds on the yields and errors to compute the bounds on key rates using Eqs. (III) and (5). The respective key rates are plotted as a function of the mean photon number ($\mu$) in Figure. (5(b)). It is established that when coincidences are taken into account, the key rate is enhanced by a substantial amount. The coincidence detection protocol helps to achieve a higher optimal mean photon number, i.e. $\mu=1.1$, as compared to using the traditional decoy-state method without coincidences, i.e. $\mu=0.8$. Comparing the optimal key rates of decoy-state protocol and EPCD protocol, we note an improvement of $\approx 64\%$ corresponding to higher achievable distances.
@@
Using the experimental setup described in the previous Section, the four protocols are executed. The protocols are systematically repeated for several signal intensities, each corresponding to distinct mean photon numbers. The experimentally acquired gain ($Q_{\mu}$) and QBER ($E_{\mu}$) are then used to estimate the optimal bounds on the secure key rate as discussed in Section. III. We have used the timestamps for recording the coincidences in addition to the single detections. The data were taken at various intensities of the source. These two cases enabled us to study four variants for our QKD protocol: Standard BB84, decoy state protocol, CD protocol, and EPCD protocol. Optimal key rates were calculated for all four cases at different values of mean photon number ($\mu$). The key rates for all the cases as a function of the mean photon number can be seen in Figure. (5(c)). It is also observed that our experiment reproduces the theoretical predictions of improvement in secure key rates. We obtained a maximum key rate of $0.17$ per pulse for the EPCD protocol.
@@
We have performed a simulation of the expected secure key rate for the four protocols as a function of distance in optical fiber. The simulated key rates are plotted in Figure. (5(d)). In this simulation we have considered the mean photon number of the signal and decoy to be $\mu=0.9$ and $\nu=0.1$, respectively. It is evident that the EPCD protocol consistently gives higher key rates than other protocols across varying distances, showcasing its effectiveness in extending communication ranges. This consistent performance highlights the EPCD protocol’s reliability and strength in addressing challenges related to communication distance.
@@ VI Conclusion
In this work, we have proposed the entrapped pulse coincidence detection (EPCD) protocol, where we have integrated two approaches to harness their collective strength. Monitoring coincidences of the signal as well as entrapped pulses allows us to detect the most practical, sophisticated PNS attacks. With an assurance of no PNS attack, we showed that including contributions from two-photon statistics, that we could achieve higher key rates.
@@ Acknowledgement
The authors acknowledge the partial funding support from DST through the QuST program. The authors are also thankful to group members of the QST lab for their valuable input, with a special thanks to Dr. Satyajeet Patil and Mr. Vardaan Mongia. RB acknowledges support from Quantum Communications Hub of the UK Engineering and Physical Sciences Research Council (EPSRC). RB thanks Mr. Lewis Wooltorton, Dr. Shashank Ranu and Prof. Roger Coleck for insightful discussions. AB thanks Prof. Marco Lucamarini and Dr. Rupesh Kumar for insightful discussions.
@@ Appendix B Analytic bounds for key-rates
$R$ $\geq\frac{1}{2}\Big{\{}-Q_{\mu}f(E_{\mu})H_{\text{bin}}(E_{\mu})$
$\qquad+Q_{1}\left(1-H_{\text{bin}}(e_{1})\right)\Big{\}},$ (21)
@@ Appendix C Numerically computing key rates
Here we present an algorithm to compute tight lower bounds for the optimization problem (III).
@@ C.1 Converting into a finite optimization problem
As the objective function of problem (III) only consists of $2$ terms with $4$ variables, removing infinite parameters from constraints suffices. Consider the following lower bound on the constraint for $Q_{\nu_{d}}$ in the optimization problem:
@@
allows us to obtain a relaxation corresponding to the constraint $Q_{\mu}e^{\mu}=\sum_{k=0}^{\infty}Y_{k}\mu^{k}/k!$ in optimization problem (III):
@@
In other words, we have formed a relaxation of the constraint given by an infinite sum of infinitely many variables through a constraint that involves a finite sum of finitely many variables. We use similar arguments to derive the following relaxations of the constraints for $E_{\mu}Q_{\mu},Q_{\nu_{i}},E_{\nu_{i}}Q_{\nu_{i}}$ in optimization problem (III):
@@
$r_{n}:=\inf$ $\Big{(}Y_{1}\mu\left(1-\Phi\left(2e_{1}-1\right)\right)$
$+Y_{2}\frac{\mu^{2}}{2}\left(1-\Phi\left((2e_{2}-1)^{2}\right)\right)\Big{)}$
@@
$\forall\lambda\in S:\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$ (26)
@@
Consider the optimization problems for $r_{n}$ and $r$. From the preceding discussion, it is evident that the constraints in problem (1) serve as relaxations for the constraints in problem (III). Consequently, the feasible set of optimization problem (III) is a subset of that of optimization problem (1). Moreover, the optimization problems designed to compute both $r$ and $r_{n}$ share identical objective functions. Thus, $r_{n}\leq r$ holds for all $n\in\mathbb{R}$.
@@ C.2 Partitioning the domain
$k_{\mathcal{P}}\leq k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ $=\tilde{f}(\mathbf{x}_{i},\mathbf{y}_{i})\leq\tilde{f}_{i}(\mathbf{x}^{*},\mathbf{y}^{*})$
$\leq f(\mathbf{x}^{*},\mathbf{y}^{*})=k.$ (30)
@@
$k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})$ $:=\min f_{i}(\mathbf{x},\mathbf{y})$
@@
$\forall j\in\{1,\cdots,M\}:h_{j}(\mathbf{x},\mathbf{y})=0$ (31)
@@
$k$ $=\min_{i}\{k_{1}(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})-\epsilon\}$
$\geq\min_{i}\{k(\mathcal{C}_{\mathbf{a}^{(i)},\mathbf{b}^{(i)}})\}-\epsilon\geq k_{\mathcal{P}}-\epsilon.$ (32)
@@ C.3 Algorithm for computing key-rates
$\xi^{\max}_{k,i}=\begin{cases}\begin{aligned} &\max\{\Phi((2a_{k}^{(i)}-1)^{k}),\\ &\Phi((2b_{k}^{(i)}-1)^{k})\}\end{aligned}&\text{if}\quad 0\notin[2a_{k}^{(i)}-1,2b_{k}^{(i)}-1]\\ \Phi(0)=1&\text{otherwise}\end{cases}$ (33)
@@
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:$
$\quad\quad Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}$
$\quad\quad\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in\{\mu,\nu_{1},\cdots,\nu_{K}\}:$
$\quad\quad E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
$\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@ Appendix D Speeding up solving the optimization problem
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
$\quad\quad\quad\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
$\quad\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@
$\forall\lambda\in S:Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
$\quad\quad\quad\leq\sum_{k=0}^{n}Y_{k}\frac{\mu^{k}}{k!}\leq Q_{\lambda}\exp(\lambda)$
$\forall\lambda\in S:E_{\lambda}Q_{\lambda}\exp(\lambda)-\Theta_{n}(\lambda)$
$\quad\quad\quad\leq\sum_{k=0}^{n}e_{k}Y_{k}\frac{\mu^{k}}{k!}\leq E_{\lambda}Q_{\lambda}\exp(\lambda)$
@@ Appendix E Rates in the presence of statistical errors
$\mathfrak{A}$ ::=:= $\left(\bigcup_{d}\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\}\right)$ (46)
$\cup\left(\bigcup_{d}\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\}\right)\cup\mathfrak{A}_{1}.$
@@
$\mathbb{P}(\mathfrak{A})$ $\leq\sum_{d}\Bigl{(}\mathbb{P}(\{|Q_{\nu_{d}}^{\mathrm{obs}}-Q_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{Q}\})$
+($\quad+\mathbb{P}(\{|E_{\nu_{d}}^{\mathrm{obs}}-E_{\nu_{d}}^{\mathrm{exp}}|\geq\delta_{\nu_{d}}^{E}\})\Bigr{)}+\mathbb{P}(\mathbb{A}_{1}).$
@@
$\mathbb{P}(\mathfrak{A})$$\leq$ $2\sum_{s=0}^{K}\left(e^{-2N_{\nu_{d}}(\delta_{\nu_{d}}^{Q})^{2}}+e^{-2N_{\nu_{i}}(\delta_{\nu_{d}}^{E})^{2}}\right)$ (49)
$+2\epsilon_{\mathrm{stat}}.$